OpenStudy (anonymous):

A graduated cylinder with a radius of 3 cm contains water to a height of 9 cm. Yann adds 15 steel balls, each with a diameter of 3 cm. What is the total volume of the contents in the cylinder after the balls are added? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (anonymous):

@Austin6i6

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

what is the volume of the cylinder with just the water in it

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

do you get it

OpenStudy (freckles):

so we need to find the volume of the water and find the 15*volume of single ball part and add that together

OpenStudy (freckles):

first lets look at the water

OpenStudy (anonymous):

ok

OpenStudy (freckles):

it has a height of 9cm and we know we put it in the cylinder which has base radius 3 we know the formula for a cylinder is h*pi*r^2

OpenStudy (freckles):

use the formula to find the volume of the water part

OpenStudy (anonymous):

so it is 254.34

OpenStudy (anonymous):

i mean thats what i got

OpenStudy (freckles):

ok now lets find the volume of one single ball

OpenStudy (freckles):

a ball is a sphere so you will want to use the volume formula for a sphere

OpenStudy (freckles):

be careful you are given the diameter of the sphere as 3cm you will need to calculate the radius of the sphere but don't worry that isn't too much of a headache because half the diameter measure is equal to the measure of the radius once you find the radius you can then move on to use that formula for the volume of a sphere

OpenStudy (anonymous):

so it will be 1.5cm

OpenStudy (freckles):

yes now plug that into the formula for the volume of a sphere so we can find the volume for one ball

OpenStudy (freckles):

\[\frac{4}{3} \cdot \pi \cdot r^3 \] this is the formula for the volume of a sphere if you do not have it in your notes

OpenStudy (anonymous):

i got 14.13cm

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

ok

OpenStudy (freckles):

great now you have 15 balls 14.13cm^3+14.13cm^3+....+14.13cm^3 pretend I wrote that number 15 times :p you can either add that number up that amount of times or you can simply do 15*(14.13) cm^3 to find the sum total of the volume of the 15 balls \[\text{ water volume } + \text{ total ball volume } \\ =254.34 \text{ cm } ^3 +15(14.13) \text{ cm}^3\]

OpenStudy (anonymous):

i got a bunch a numbers when i multiplyed 14.13 15 times

OpenStudy (freckles):

15(14.13)= only a 5 night number

OpenStudy (freckles):

digit*

OpenStudy (freckles):

not night lol

OpenStudy (anonymous):

lol is it 211.95

OpenStudy (freckles):

yes now add that result to 254.34

OpenStudy (anonymous):

so 466.29

OpenStudy (freckles):

yeah

OpenStudy (freckles):

cm^3

OpenStudy (anonymous):

thats the answer

OpenStudy (freckles):

\[466.29 cm^3 \text{ is the total volume }\]

OpenStudy (anonymous):

thank you so much

OpenStudy (freckles):

np

OpenStudy (anonymous):

can you help me with 2 more